# On the Brun–Titchmarsh theorem

### v1

@@ 5.2. Approaching by multiplicative characters
$R(D)$ $=\frac{1}{\varphi(q)}\frac{1}{2\pi}\int_{\mathbf{R}}\widetilde{\varPhi}(\tfrac{1}{2}+it)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(\tfrac{1}{2}+it,\chi)L(\tfrac{1}{2}+it,\chi)\mathrm{d}t+O(1),$
@@
$\mathcal{D}(s,\chi)=\sum_{d\leqslant D}\lambda(d)\chi(d)d^{-s}.$
@@
$\int_{\mathbf{R}}|\widetilde{\varPhi}(\tfrac{1}{2}+it)|(1+|t|)^{3}\mathrm{d}t$ $\ll x^{\frac{1}{2}}(\log x)^{12},$
@@
(5.5) $\Big{|}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(s,\chi)L(s,\chi)\Big{|}\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
Due to the well-factorization of $\lambda$, we may write
[…]
where $\boldsymbol{\alpha}$, $-β\bbeta$ and $\boldsymbol{\gamma}$ are of level $M$, $N$ and $W$ respectively, and
@@
$\mathcal{M}(s,\chi)=\sum_{m\sim M}\alpha_{m}\chi(m)m^{-s},\ \ \mathcal{N}(s,\chi)=\sum_{n\sim N}\beta_{n}\chi(n)n^{-s},\ \ \mathcal{W}(s,\chi)=\sum_{w\sim W}\gamma_{w}\chi(w)w^{-s}.$
@@
(5.7) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)L(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}$
for $\Re(s)=\frac{1}{2}.$ Recall that $\varPhi$ is supported in $[y,x+y]$. For $m\sim M,n\sim N,w\sim W,$ it is clear that $\varPhi(mnwl)$ vanishes unless $l\asymp x/(MNW)$. As argued above, an alternative approach to (5.4) is proving
(5.8) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036},$
@@
$\mathcal{L}(s,\chi)=\sum_{l\in\mathbf{Z}}\phi\Big{(}\frac{l}{L}\Big{)}\chi(l)l^{-s},\ \ \ L=\frac{x}{MNW}.$
@@
We assume that $W=1$ in (5.6), so that (5.8) reduces to
(5.9) $\Xi:=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
By Cauchy–Schwarz inequality, we have
$\Xi^{2}\ll(M+q)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|^{2}.$
From Lemma 3.4 it follows that
$\Xi^{2}\ll|s|^{\frac{9}{2}}(M+q)N^{\frac{1}{2}}q^{\frac{\vartheta+1}{2}+\varepsilon}(1+Nq^{-\frac{1}{2}})^{\frac{3}{2}}+(M+q)(|s|^{\frac{9}{2}}q+xM^{-1}(\log q)^{-A})$
for any $A>0$, provided $NL>q^{1+\varepsilon}.$ It is easy to check that (5.9) is valid as long as
$M=x/q^{1+\varepsilon},\ \ N=q^{\frac{5-2\vartheta}{8}-\varepsilon}$
if $q<x^{\frac{1}{2}-}$, in which case we may take $D=x/q^{\frac{3+2\vartheta}{8}+2\varepsilon}$ in Proposition 5.1. This proves the first part of Theorem 1.1 readily.
We now assume that $q$ is a large prime. As argued above, (5.7) reduces to
(5.10) $\Theta:=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
@@
(5.11) $\Theta^{4}\ll\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)|^{\frac{4}{3}}\Big{)}^{3}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\Big{)}.$
If $N^{2}\leqslant q^{\frac{1}{4}}$, we infer from Lemma 3.5 that
(5.12) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\ll|s|^{12}q^{1+\varepsilon}.$
To bound the $\frac{4}{3}$-th moment of $\mathcal{M}(s,\chi),$ we assume that $\boldsymbol{\alpha}=\boldsymbol{\alpha}_{1}*\boldsymbol{\alpha}_{2}*\boldsymbol{\alpha}_{3}$ is a triple convolution of three coefficients $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}$ and $\boldsymbol{\alpha}_{3},$ each of which is of level $M^{\frac{1}{3}}$. This is reasonable since $\boldsymbol{\alpha}$ comes from the well-factorable weight constructed by Iwaniec [Iw82]; see Lemma 2.1. Note that $\mathcal{M}_{j}(s,\chi)$ denotes the Dirichlet polynomial generated by $\boldsymbol{\alpha}_{j}$. We now derive from Lemma 3.1 that
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)|^{\frac{4}{3}}$ $=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}_{1}(s,\chi)\mathcal{M}_{2}(s,\chi)\mathcal{M}_{3}(s,\chi)|^{\frac{4}{3}}$
$\ll\prod_{1\leqslant j\leqslant 3}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}_{j}(s,\chi)|^{4}\Big{)}^{\frac{1}{3}}$
$\ll(M^{\frac{2}{3}}M^{\varepsilon}+q).$
from which and (5.12), (5.11), we find (5.10) is valid as long as
$M=x/q^{\frac{1}{2}+2\varepsilon},\ \ N=q^{\frac{1}{8}},$
if $q<x^{\frac{1}{2}}(\log x)^{-2036}$, in which case we may take $D=x/q^{\frac{3}{8}+2\varepsilon}$ in Proposition 5.1. The second part of Theorem 1.1 is established.
By the way, if assuming Conjecture 1.1, we may take $N=q^{\delta/2}$, so that the choice $D=x/q^{\frac{1-\delta}{2}+2\varepsilon}$ is admissible in Proposition 5.1. This proves Theorem 1.7.

### v3

@@ 5.1. Approaching by multiplicative characters
$R^{+}(D)$ $=\frac{1}{\varphi(q)}\frac{1}{2\pi}\int_{\mathbf{R}}\widetilde{\varPhi}(\tfrac{1}{2}+it)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(\tfrac{1}{2}+it,\chi)L(\tfrac{1}{2}+it,\chi)\mathrm{d}t+O(1),$
@@
$\mathcal{D}(s,\chi)=\sum_{d\leqslant D}\lambda^{+}(d)\chi(d)d^{-s}.$
@@
$\int_{\mathbf{R}}|\widetilde{\varPhi}(\tfrac{1}{2}+it)|(1+|t|)^{4}\mathrm{d}t$ $\ll x^{\frac{1}{2}}(\log x)^{12},$
@@
(5.3) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(s,\chi)L(s,\chi)\ll|s|^{4}x^{\frac{1}{2}}(\log x)^{-2036}.$
Due to the well-factorization of $\lambda^{+}$, we write
(5.4) $-β\displaystyle\lambda^{+}=\boldsymbol{\alpha}*\bbeta,$
where $\boldsymbol{\alpha}$ and $-β\bbeta$ are of level $M$, $N$, respectively, and
@@
$\mathcal{M}(s,\chi)$ $=\sum_{m\sim M}\alpha_{m}\chi(m)m^{-s},$
$\mathcal{N}(s,\chi)$ $=\sum_{n\sim N}\beta_{n}\chi(n)n^{-s},$
$\mathcal{P}(s,\chi)$ $=\sum_{p\sim P}\chi(p)p^{-s}.$
@@
(5.5) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|\ll|s|^{4}x^{\frac{1}{2}}(\log x)^{-2036}$
for $\Re(s)=\frac{1}{2}.$ Recall that $\varPhi$ is supported in $[y,x+y]$. For $m\sim M,n\sim N,$ it is clear that $\varPhi(mnl)$ vanishes unless $l\asymp x/(MN)$. As argued above, an alternative approach to (5.1) is proving
(5.6) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{4}x^{\frac{1}{2}}(\log x)^{-2036},$
@@
$\mathcal{L}(s,\chi)=\sum_{l\in\mathbf{Z}}\phi\Big{(}\frac{l}{L}\Big{)}\chi(l)l^{-s},\ \ \ L=\frac{x}{MN}.$
@@
In a similar way, the proof (5.2) can be reduced to proving that
(5.7) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{P}(s,\chi)\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|\ll|s|^{4}x^{\frac{1}{2}}(\log x)^{-2036}$
or
(5.8) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{P}(s,\chi)\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{4}x^{\frac{1}{2}}(\log x)^{-2036},$
where $\Re(s)=\frac{1}{2}.$
@@
We would like to estimate the error terms in Proposition 4.2. They correspond to the desired inequalities in (5.1) and (5.2). The first one will be dealt with using Lemma 2.5, and in the treatment to the second one we utilize the convolution structures of the coefficients in Dirichlet polynomials, so that Lemma 2.6 applies. Assume that $x^{\frac{9}{20}}\leqslant q<x^{\frac{1}{2}-}$ since the case $q<x^{\frac{9}{20}}$ has been dealt with by Iwaniec [Iw82].
Denote by $\Xi_{1}$ the LHS of (5.6). By Cauchy–Schwarz inequality and Lemma 2.1, we have
$\Xi_{1}^{2}\ll(M+q)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|^{2},$
where $MN=D$ and $M,N\geqslant 1.$ From Theorem 2.1, it follows that
$\Xi_{1}^{2}\ll|s|^{\frac{7}{2}}(M+q)(q+N^{\frac{33}{20}}q^{\frac{3}{20}}+N^{\frac{15}{8}})q^{\varepsilon}.$
Note that the choice
(5.9) $M=xq^{-1-\varepsilon},\ \ N=q^{\frac{17}{33}-\varepsilon}$
is valid to guarantee the bound (5.6), from which we obtain a level $D=xq^{-\frac{16}{33}-2\varepsilon}$ in (5.1).
If appealing to Proposition 4.1 directly, we may choose $C(\varpi)=66/(33-16\varpi)$ in (1.3), i.e., one can take $C^{*}(\varpi)=0$ in (1.4). The positive choices of $C^{*}(\varpi)$ in (1.5) and (1.6) are due to the task of Proposition 4.2, in which the trilinear structure in the error term becomes very crucial.
[…]
$\sum_{p\sim P}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)$ $\geqslant\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}A(pd)$
$=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)+\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(p)\\ \omega(d)>(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)=\pi_{1}^{*}+\pi_{2}^{*},$
say. Recall that $A(d)$ is defined by (4.6). Hence
$\pi_{2}^{*}$ $\ll\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ \omega(d)>(\log x)^{1/3}\end{subarray}}|\lambda_{d}^{-}|\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ pd\mid n\end{subarray}}\varPhi(n)$
$\ll 2^{-(\log x)^{1/3}}\sum_{d\leqslant 2PD^{*}}2^{\omega(d)}\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ d\mid n\end{subarray}}\varPhi(n)$
$\ll 2^{-(\log x)^{1/3}}\sum_{\begin{subarray}{c}n\leqslant 2x\\ n\equiv a\,(\operatorname{mod}{q})\end{subarray}}3^{\omega(n)}.$
This is at most $O(x/(q\log^{3}x))$ by Shiu [Sh80, Theorem 2], provided that $D>x^{0.01}$ and $q<x^{0.99},$ in which case we have
$\sum_{p\sim P}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)$ $\geqslant\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\end{subarray}}\lambda_{d}^{-}A(pd)$
$=\sum_{p\sim P}\sum_{\begin{subarray}{c}d\leqslant D^{*}\\ d\mid P(p)\\ \omega(d)\leqslant(\log x)^{1/3}\end{subarray}}\lambda_{d}^{-}A(pd)+O\Big{(}\frac{x}{q(\log x)^{3}}\Big{)}.$
Hence it suffices to consider those $\lambda_{d}^{-}$ supported on $\{d\leqslant D^{*}:\mu^{2}(d)=1,\omega(d)\leqslant(\log x)^{1/3}\}.$ With such modifications to the definitions of $\mathcal{M}(s,\chi)$ and $\mathcal{N}(s,\chi)$, it follows from Lemma 2.6 that
$\Xi_{2}^{2}$ $\ll|s|^{\frac{9}{2}}(M+q)P^{\frac{1}{2}}N^{2}q^{\frac{\vartheta+1}{2}}(Pq^{-\frac{1}{2}}+1)q^{\varepsilon}+|s|(M+q)q^{1+\varepsilon}+|s|^{\frac{9}{2}}(M+q)Pq^{\frac{1}{2}+\varepsilon}$
for any $\varepsilon>0,$ where $MN=D^{*}$ and $M,N\geqslant 1.$ To guarantee the bound (5.6), we hope
$MN^{2}$ $\leqslant x^{1-\varepsilon}q^{-\frac{\vartheta+1}{2}}P^{-\frac{1}{2}}(Pq^{-\frac{1}{2}}+1)^{-1},$
$N^{2}$ $\leqslant x^{1-\varepsilon}q^{-\frac{\vartheta+3}{2}}P^{-\frac{1}{2}}(Pq^{-\frac{1}{2}}+1)^{-1},$
$M$ $\leqslant x^{1-\varepsilon}q^{-1},$
$M$ $\leqslant x^{1-\varepsilon}q^{-\frac{1}{2}}P^{-1}.$
The above inequalities also require an additional condition $P\leqslant x^{1-\varepsilon}q^{-\frac{3}{2}}$, which we henceforth assume. In particular, we may take
$M=x^{1-\varepsilon}q^{-1}\min\{1,P^{-1}q^{\frac{1}{2}}\},\ \ N=x^{-\varepsilon}q^{\frac{1-\vartheta}{4}}P^{-\frac{1}{4}}.$
Hence (5.2) holds with
$D^{*}=x^{1-2\varepsilon}q^{-\frac{3+\vartheta}{4}}P^{-\frac{1}{4}}\min\{1,P^{-1}q^{\frac{1}{2}}\},$
so that one may take
$D_{p}=x^{1-2\varepsilon}q^{-\frac{3+\vartheta}{4}}p^{-\frac{1}{4}}\min\{1,p^{-1}q^{\frac{1}{2}}\}$
in Proposition 4.2. This gives
$\pi(x;q,a)$ $\leqslant(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{x}{\varphi(q)}\Big{\{}\frac{1}{\log w}F\Big{(}\frac{\log(xq^{-\frac{16}{33}})}{\log w}\Big{)}$
$\ \ \ \ \ -\sum_{w\leqslant p<z}\frac{1}{p\log p}f\Big{(}\frac{\log(xq^{-\frac{3+\vartheta}{4}}p^{-\frac{1}{4}}\min\{1,p^{-1}q^{\frac{1}{2}}\})}{\log p}\Big{)}\Big{\}}$
for $z\leqslant x^{1-\varepsilon}q^{-\frac{3}{2}}.$
We would like to specialize $z,w$ such that the argument in the above $F(\cdot)$ falls into $[3,5]$ and that in $f(\cdot)$ falls into $[2,4].$ To do so, we put
$z=x^{\kappa_{1}-\varepsilon},\ \ w=x^{\kappa_{2}},\ \ \kappa_{1}=\min\Big{\{}1-\frac{3}{2}\varpi,~{}\frac{8-(1+2\vartheta)\varpi}{28}\Big{\}},\ \ \kappa_{2}=\frac{1}{5}\Big{(}1-\frac{16}{33}\varpi\Big{)}.$
By the Prime Number Theorem and partial summation,
$\pi(x;q,a)$ $\leqslant(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{x}{\varphi(q)}\Big{\{}\frac{1}{\log w}F\Big{(}\frac{\log(xq^{-\frac{16}{33}})}{\log w}\Big{)}$
$\ \ \ \ \ -\int_{w}^{z}\frac{1}{t(\log t)^{2}}f\Big{(}\frac{\log(xq^{-\frac{3+\vartheta}{4}}t^{-\frac{1}{4}}\min\{1,t^{-1}q^{\frac{1}{2}}\})}{\log t}\Big{)}\mathrm{d}t\Big{\}}$
$=(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{c\cdot x}{\varphi(q)\log x},$
where
$c$ $=\frac{1}{\kappa_{2}}F\Big{(}\frac{1-\frac{16}{33}\varpi}{\kappa_{2}}\Big{)}-\int_{\kappa_{2}}^{\varpi/2}f\Big{(}\frac{4-(3+\vartheta)\varpi)}{4u}-\frac{1}{4}\Big{)}\frac{\mathrm{d}u}{u^{2}}-\int_{\varpi/2}^{\kappa_{1}}f\Big{(}\frac{4-(1+\vartheta)\varpi}{4u}-\frac{5}{4}\Big{)}\frac{\mathrm{d}u}{u^{2}}.$
Recalling the explicit expressions (4.2) and (4.3) for $F$ and $f,$ we write
$c$ $=\frac{2\mathrm{e}^{\gamma}}{1-\frac{16}{33}\varpi}\Big{(}1+\int_{2}^{4}\frac{\log(t-1)}{t}\mathrm{d}t\Big{)}-\frac{4\cdot 2\mathrm{e}^{\gamma}}{4-(3+\vartheta)\varpi}\int_{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}^{\frac{5(4-(3+\vartheta)\varpi)}{4(1-\frac{16}{33}\varpi)}-\frac{1}{4}}\frac{\log(t-1)}{t}\mathrm{d}t$
$\ \ \ \ -\frac{4\cdot 2\mathrm{e}^{\gamma}}{4-(1+\vartheta)\varpi}\int_{\max\{\frac{4-(1+\vartheta)\varpi}{2(2-3\varpi)}-\frac{5}{4},~{}2\}}^{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}\frac{\log(t-1)}{t}\mathrm{d}t.$
Hence we may take
$C(\varpi)$ $=\frac{66}{33-16\varpi}-C^{*}(\varpi)$
in (1.3), where
$C^{*}(\varpi)$ $=-\frac{66}{33-16\varpi}\int_{2}^{4}\frac{\log(t-1)}{t}\mathrm{d}t+\frac{8}{4-(3+\vartheta)\varpi}\int_{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}^{\frac{165(4-(3+\vartheta)\varpi)}{4(33-16\varpi)}-\frac{1}{4}}\frac{\log(t-1)}{t}\mathrm{d}t$
$\ \ \ \ +\frac{8}{4-(1+\vartheta)\varpi}\int_{\max\{\frac{4-(1+\vartheta)\varpi}{2(2-3\varpi)}-\frac{5}{4},~{}2\}}^{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}\frac{\log(t-1)}{t}\mathrm{d}t.$
This proves the first part of Theorem 1.1 readily.
We now assume that $q$ is a large prime, and denote by $\Theta$ the LHS of (5.7), i.e.,
$\Theta=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{P}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|.$
@@
(5.10) $\Theta^{4}\ll\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)|^{2}\Big{)}^{2}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{P}(s,\chi)|^{4}\Big{)}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\Big{)}.$
From Lemma 2.7 we infer
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\ll|s|^{16}(N^{4}q^{-\frac{1}{2}}+1)q^{1+\varepsilon}.$
Together with Lemma 5.1, we find
$\Theta\ll|s|^{4}(M+q)^{\frac{1}{2}}(P^{2}+q)^{\frac{1}{4}}(Nq^{-\frac{1}{8}}+1)q^{\frac{1}{4}+\varepsilon}.$
Hence (5.7) is valid as long as
$M=x^{1-\varepsilon}q^{-1}\min\{1,P^{-1}q^{\frac{1}{2}}\},\ \ N=q^{\frac{1}{8}},$
if $q<x^{\frac{1}{2}-}$, $P\leqslant x^{1-\varepsilon}q^{-\frac{3}{2}}$. Upon the above choice, we may take
$D_{p}=x^{1-\varepsilon}q^{-\frac{7}{8}}\min\{1,p^{-1}q^{\frac{1}{2}}\}$
in Proposition 4.2.
As argued above, we arrive at
$\pi(x;q,a)$ $\leqslant(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{x}{\varphi(q)}\Big{\{}\frac{1}{\log w}F\Big{(}\frac{\log(xq^{-\frac{16}{33}})}{\log w}\Big{)}$
$\ \ \ \ \ -\int_{w}^{z}\frac{1}{t(\log t)^{2}}f\Big{(}\frac{\log(xq^{-\frac{7}{8}}\min\{1,t^{-1}q^{\frac{1}{2}}\})}{\log t}\Big{)}\mathrm{d}t\Big{\}}$
with
$z=x^{\min\{1-\frac{3}{2}\varpi,~{}\frac{1}{3}-\frac{1}{8}\varpi\}-\varepsilon},\ \ \ w=x^{\frac{1}{5}(1-\frac{16}{33}\varpi)}.$
Hence
$\pi(x;q,a)$ $\leqslant(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{x}{\varphi(q)\log x}\Big{\{}\frac{5F(5)}{1-\frac{16}{33}\varpi}-\int_{\frac{1}{5}(1-\frac{16}{33}\varpi)}^{\frac{1}{2}\varpi}f\Big{(}\frac{1-\frac{7}{8}\varpi}{u}\Big{)}\frac{\mathrm{d}u}{u^{2}}$
$\ \ \ \ -\int_{\frac{1}{2}\varpi}^{\min\{1-\frac{3}{2}\varpi,~{}\frac{1}{3}-\frac{1}{8}\varpi\}}f\Big{(}\frac{1-\frac{3}{8}\varpi}{u}-1\Big{)}\frac{\mathrm{d}u}{u^{2}}\Big{\}}.$
This establishes the second part of Theorem 1.1 by substituting the explicit expressions (4.2) and (4.3) for $F$ and $f.$